OpenStudy (anonymous):

find the area of the region bounded by the given curves. y = 1-2x^2, y = |x|

ganeshie8 (ganeshie8):

what have you tried ? where are you stuck at ? show your attempt

OpenStudy (anonymous):

well i have tried equaling the equations to each other to find the intercepts

OpenStudy (anonymous):

but im stuck there

ganeshie8 (ganeshie8):

thats a good start

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

you can split y = |x| into a piecewise function : \[|x| = \left\{\begin{array}{r} -x&:&x\lt0\\x&:&x\ge 0 \end{array}\right.\]

OpenStudy (anonymous):

k

ganeshie8 (ganeshie8):

for the left side intersection point, you solve below : \[ y = 1-2x^2\] \[y = -x\]

OpenStudy (anonymous):

how did u get -x?

OpenStudy (anonymous):

he got the -x from the absolute value function

ganeshie8 (ganeshie8):

^

OpenStudy (anonymous):

k

OpenStudy (anonymous):

idk how to solve it. thats where im stuck

ganeshie8 (ganeshie8):

set them equal, you will get a quadratic

ganeshie8 (ganeshie8):

\[-x = 1-2x^2\] \[2x^2-x-1 =0\] can you solve x ?